# Lie symmetry analysis of the two-Higgs-doublet model field equations

### v1

@@ 4.3.3 The inequivalent Lie point symmetries of the 2HDM
All the 2-dimensional Lie algebras found in the analysis are equivalent, since a rotation
@@
of $\mathfrak{su}(2)_{\text{HF}}$ generators will be an inner automorphism of $\mathfrak{su}(2)_{\text{HF}}$, which means it may be implemented by a matrix conjugation by a matrix $U\in\mathsf{SU}(2)$, i.e. $UH_{1}U^{\dagger}=H_{2}$, etc. Moreover, a conjugation by a $\mathsf{SU}(2)$ matrix $U$ corresponds to a change of Higgs doublet basis, which means all the 2-dimensional symmetries we found really are the same symmetry transformation, expressed in different Higgs bases. Here the other element $X_{Y}$ of the 2-dimensional algebras corresponds to a complex phase, and commutes with all other (series of) generators, including $U$, and is not affected by the automorphism. Finally, the rotation of generators (4.69) is an automorphism since it conserves the commutator, and it is inner because all automorphisms of $\mathfrak{su}(2)$ are inner (its Dynkin diagram is just a dot with no non-trivial diagram automorphisms). Hence we have shown there are only three possible, inequivalent scalar Lie point symmetry algebras of the 2HDM, namely
@@
and they are all strictly variational due to Theorem 1, consistent with the results in [21, 32]. Gauge symmetries are not included in (4.70), since we only have considered (purely) scalar symmetries. We have in addition demonstrated something new, namely that there are no scalar divergence or scalar non-variational Lie point symmetries in the 2HDM, since Lie’s method finds all symmetries of systems of PDEs, cf. (2.17). The absence of divergence and non-variational Lie point symmetries in the 2HDM stands in contrast to the situation in singlet extensions of the SM, where such symmetries do appear [50].
Furthermore, we have tested the soundness of our implementation of Lie symmetry analysis for models with many parameters, variables and reparametrization freedom, by applying the method on a well-understood example and replicating well-known results. Of course, the derivation in [21] of the strict variational point symmetries of the 2HDM is more effective, transparent and elegant, and also includes discrete symmetries in the same proof. However, the method applied in this section has the advantage that it is universal, and that it detects all divergence and non-variational symmetries as well as strict variational symmetries. Lie symmetry analysis could, at least in principle, be applied to any model, including any NHDM, in contrast to methods only considering the possible symmetry transformations of gauge invariant, scalar bilinears, cf. (4.9). The method applied in this work also yields the exact parameter conditions for all symmetries, although this may not be an essential feature in a classification of symmetries.
@@ 5 Summary and outlook
We have demonstrated, by example, how Lie symmetry analysis of field equations can be applied to quantum physical models with many variables, parameters and reparametrization freedom. In our analysis of the two-Higgs-doublet model (2HDM), we reproduced the well-known strict variational Lie point symmetries of the model and also showed the model forbids divergence and non-variational Lie point symmetries. This observation is essential: Divergence symmetries are, under the usual boundary conditions, symmetries of the action and will, provided the measure of Feynman’s path integral is invariant, be preserved under quantization in the same manner as strict variational symmetries. Moreover, we proved three general results that may simplify Lie symmetry analysis for a wide class of particle physics models. The most relevant result for the present work was Theorem 1, which enables us to decide whether scalar Lie point symmetries are variational for a large class of theories, simply by considering the actions of the symmetries on the kinetic terms. This criterion is independent of the possibly numerous parameter conditions accompanying the different potentials that may realize the same symmetry. The two other results, Proposition 2.5 and Corollary 1, were not strictly necessary for the present analysis, but they may substantially simplify the computations of variational, scalar Lie point symmetries of other, larger models of the type NHDM+KS. This is done by reducing the spacetime dimension, and hence the number of spacetime variables and gauge fields, while getting exactly the same results for scalar variational symmetries. Lie symmetry analysis only detects continuous (Lie) symmetries. Nevertheless, any discrete symmetries not detected by Lie’s method, may be identified through considering the automorphism groups of the Lie symmetry algebras computed through Lie symmetry analysis. In this way, the maximal, faithfully acting variational symmetry group of a theory may be determined. We leave a systematic study of such automorphisms and their role in detecting discrete symmetries of large quantum physical models for future work.
@@ Acknowledgments
The most general NHDM+KS Lagrangian in $d$ spacetime dimensions was given by (2.50). We may expand the Lagrangian (2.50) in purely scalar terms and terms linear and quadratic in the gauge fields, as follows
@@
represent the gauge bosons and where e.g. $\mu\in\{0,1,2,3\}$ for $\mathcal{L}_{4}$ while $\mu=0$ for $\mathcal{L}_{1}$. Moreover, $T_{\text{GB}}$ is the sum of the kinetic terms of the gauge bosons given by (2.51), which will be irrelevant for the proof of Proposition 2.5, since we only will be considering scalar symmetries. The relevant 1-prolongation of the scalar infinitesimal generator for the $d$-dimensional case is given by
@@
First we consider strict variational symmetries, i.e. symmetries with $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$ (no sum over $d$), cf. (2.21) with $\xi=0$.
Denote $\partial_{\phi_{i}}\Phi_{j}\equiv\Phi_{j,i}$ and similar for Hermitian conjugated fields. Then e.g. $\Phi_{1,4}^{\dagger}=(0,-i)$. The effect of $\operatorname{pr}^{(1)}X_{d}$ of (A.3) on $\mathcal{L}_{d}$, given by (A), may be written
@@
where $0\leq\mu\leq d-1$ and $1\leq K$. Assume that $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}=0$ holds, we would like to show that then $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$ holds, with the same $\eta$’s. This will mean that a scalar symmetry of the $d=1$ theory also is a scalar symmetry of the theory with arbitrary dimension $d$. Now, $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$, since all other terms are proportional to gauge bosons $G^{\mu}$ or derivatives $\partial\phi_{j}/\partial x_{\mu}$ (which includes terms $\partial_{\mu}s_{m}$), and hence cannot cancel $\eta_{i}{\partial_{\phi_{i}}V(\Phi,s)}{}$. This infers that $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$ in $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ as well, since it is the same expression, when suppressing spacetime variables. The other terms all depend on $\mu$, and since they cancel for $\mu=0$ in $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}$, they also have to cancel for each $\mu$ in $\operatorname{pr}^{(1)}X_{4}\mathcal{L}_{4}$, since the structure of the terms is the same for all $\mu$.
@@
Conversely, assume $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$, that is, we have a symmetry of the theory of some fixed, but arbitrary dimension $d>1$. Then (A) holds where $\mu$ this time is a sum from 0 to d1d-1. We will show that then $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}=0$ for the same $\eta$’s, which means the $d=1$ theory has the same symmetry. Again, since all other terms are proportional to gauge bosons or derivatives, $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$, and the same is true in the $d=1$ theory. Moreover, the other terms have to cancel for each $\mu$ individually: This is so because terms involving gauge bosons $G^{\mu}$ cannot be canceled by terms involving gauge bosons $G^{\nu}$ for $\nu\neq\mu$, since they are independent fields. Moreover the term $\frac{\partial\phi_{j}}{\partial{x_{\mu}}}\frac{\partial\eta_{i}}{\partial\phi_{j}}\frac{\partial\phi_{i}}{\partial x^{\mu}}$ for one specific, fixed $\mu$ cannot be cancelled by a term of the same type with fixed spacetime index $\nu\neq\mu$, since $\eta_{i}$ doesn’t involve spacetime derivatives of fields. There are no other terms that can cancel these terms either. So terms containing $\mu$ in (A) cancel for each choice of $\mu$ separately, and hence they will cancel for $\mu=0$ in the $d=1$ theory as well, and the symmetry for the $d>1$ theory is also a symmetry for the $d=1$ theory. So the scalar, SVSs for $\mathcal{L}_{d}$ and $\mathcal{L}_{1}$ are the same, hence they are the same for $\mathcal{L}_{d}$ for all $d\in\mathbb{N}$.
@@
We now turn to the case where $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=\operatorname{Div}\beta\equiv d_{\mu}\beta^{\mu}$ for some local vector field $\beta^{\mu}(x,\phi,\phi_{,\alpha},\phi_{,\alpha\beta})$, that is not identically zero. The first order infinitesimal variation $\delta\mathcal{L}_{d}=\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ is a divergence if and only if the Euler operator $E$ applied on $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ is zero, cf. (2.26), which means
@@
for all $1\leq j\leq(4N+K+4d)$, where the dependent variables $y^{j}$ are the $4N+K$ Higgs field components and the $4d$ gauge field components, and where $d$ again is the number of independent variables (the dimension). First, assume $E(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$, i.e. we have a divergence symmetry in the $d=1$ theory. The term $\eta_{i}{\partial_{\phi_{i}}V(\Phi,s)}{}$ only depends on undifferentiated scalars, hence if $E_{j}$ corresponds to a Higgs field $y^{j}=\phi_{j}$, then $E_{j}(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$ infer $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$. This is so because the term $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))$ cannot be canceled by any other term in $E_{j}(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})$ since they all will be proportional to either terms of the type $G_{\mu}$, $\partial_{\mu}\phi_{k}$ or a $\partial_{\mu}\partial^{\mu}\phi_{k}$ (all with $\mu=0$), in contrast to the term $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))$. Hence $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$ also will be true in the $d>1$ theory, since these terms are the same for all $d$, suppressing spacetime variables. All the remaining terms of (A) have got two factors of the type $G_{\mu}$ or $\partial_{\mu}\phi_{k}$ for $\mu=0$, and no matter which $E_{i}$ we apply, at least one such factor survives in each non-annihilated term, so all these terms are tagged with the index $\mu=0$. Since the $d>1$ dimensional theory have $d$ copies of these terms, where each copy is tagged with an index $\mu$, $0\leq\mu\leq d-1$, each of the $d$ copies have to vanish separately, since they are structurally the same, only distinguished by the index $\mu$. Hence we also have $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=0$.
@@
Conversely, assume $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=0$, i.e. we have a divergence symmetry in $d>1$ spacetime dimensions. By the same reason as before the terms $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$ in the $d>1$ theory, and hence the same is true in the $d=1$ theory, since those terms are the same for all choices of $d$. The remaining terms of $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})$ can be grouped in $d$ distinct, non-overlapping categories determined by the value of $\mu$ in factors of the types $G_{\mu}$, $\partial_{\mu}\phi_{k}$ or $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$. The terms in each category have to cancel separately, since all terms in one category are linearly independent of any other term in another category. [^12] Note that the the conditions for variational symmetries, including divergence symmetries, don’t presuppose that the Euler-Lagrange equations of the Lagrangian holds. Hence there is no relation between $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$ and $\partial_{\check{\nu}}\partial^{\check{\nu}}\phi_{k}$ for $\check{\mu}\neq\check{\nu}$. Hence the terms in the category $\mu=0$ cancel, and hence $E(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$. Therefore, the scalar symmetries of $\mathcal{L}_{d}$ are the same for all $d$, since $\mathcal{L}_{d}$ have the same scalar symmetries as $\mathcal{L}_{1}$ for all $d$. Finally, we note that the arguments hold equally well if $K=0$, i.e. we have a pure NHDM. ∎
@@
We don’t expect Proposition 2.5 to hold in general for non-variational symmetries, since the kinetic terms of the gauge bosons consist of cross terms of different spacetime derivatives, and these terms will now matter, since a symmetry of the Euler-Lagrange equations here infer that
@@
and the kinetic terms of the gauge bosons will be involved in the ”on-shell” condition $E(\mathcal{L}_{d})=0$, while this was not the case for the ”off-shell” (i.e. variational), scalar symmetries in Proposition 2.5. In contrast, the condition for a variational symmetry here is
@@
with no assumption on the validity of the Euler-Lagrange equations. Moreover, for $d=1$ the kinetic terms of the gauge bosons are zero due to the antisymmetry of the spacetime indices, while they are non-zero for $d>1$, which induces an asymmetry between the $d=1$ and $d>1$ cases in the on-shell condition $E(\mathcal{L}_{d})=0$ for the two cases. Finally, the Euler-Lagrange equations will infer dependencies between derivatives like $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$ and $\partial_{\check{\nu}}\partial^{\check{\nu}}\phi_{k}$ for $\check{\mu}\neq\check{\nu}$, in contrast to the off-shell situation in Proposition 2.5, cf. footnote 12.

### v2

@@ 4.3.3 The inequivalent Lie point symmetries of the 2HDM
All the 2-dimensional Lie algebras found in the analysis are equivalent, because a rotation
@@
of $\mathfrak{su}(2)_{\text{HF}}$ generators will be an inner automorphism of $\mathfrak{su}(2)_{\text{HF}}$, which means it may be implemented by a matrix conjugation by a matrix $U\in\mathsf{SU}(2)$, that is, $UH_{1}U^{\dagger}=H_{2}$, etc. Moreover, a conjugation by a $\mathsf{SU}(2)$ matrix $U$ corresponds to a change of Higgs doublet basis, which means all the 2-dimensional symmetries we found really are the same symmetry transformation, expressed in different Higgs bases. Here the other element $X_{Y}$ of the 2-dimensional algebras corresponds to a complex phase, and commutes with all other (series of) generators, including $U$, and is not affected by the automorphism. Finally, the rotation of generators (4.69) is an automorphism because it conserves the commutator, and it is inner because all automorphisms of $\mathfrak{su}(2)$ are inner (its Dynkin diagram is just a dot with no non-trivial diagram automorphisms). Hence, we have shown that there are only three possible, inequivalent scalar Lie point symmetry algebras of the 2HDM, namely
@@
and they are all strictly variational according to Theorem 1, consistent with the results in [21] and [32]. Gauge symmetries are not included in (4.70) because we considered only (purely) scalar symmetries. In addition, we have demonstrated something new, namely, that there are no scalar divergence or scalar non-variational Lie point symmetries in the 2HDM, as Lie’s method finds all symmetries of systems of PDEs, cf. (2.19). The absence of divergence and non-variational Lie point symmetries in the 2HDM stands in contrast to the situation in singlet extensions of the SM, where such symmetries do appear [49].
Furthermore, we have tested the soundness of our implementation of Lie symmetry analysis for models with many parameters, variables and reparametrization freedom by applying the method to a well-understood example and replicating well-known results. Of course, the derivation in [21] of the strict variational point symmetries of the 2HDM is more effective, transparent and elegant, and includes discrete symmetries in the same proof. However, the method applied in this section has the advantage that it is universal and detects all divergence and non-variational symmetries as well as strict variational symmetries. Lie symmetry analysis could, at least in principle, be applied to any model, including any NHDM, in contrast to methods that only consider the possible symmetry transformations of gauge invariant scalar bilinears, cf. (4.9). The method applied in this study also yields the exact parameter conditions for all symmetries, although this may not be an essential feature in a classification of symmetries.
@@ 5 Summary and outlook
We have demonstrated, by example, how Lie symmetry analysis of field equations can be applied to particle physics models with many variables, parameters, and reparametrization freedom. In our analysis of the two-Higgs-doublet model (2HDM), we reproduced its well-known strict variational Lie point symmetries and, for the first time, showed that the model admits neither divergence nor non-variational Lie point symmetries. Hence, the maximal realizable Lie point symmetry algebras of the 2HDM are generated exhaustively by the strict variational symmetries. This observation is essential: Under the usual boundary conditions, divergence symmetries are symmetries of the action and, provided that the path-integral measure is invariant, they should be preserved under quantization in the same way as strict variational symmetries.
Moreover, we proved three general results that can simplify Lie symmetry analysis for a wide class of particle physics models. The most relevant result for the present work is Theorem 1, which enables us to decide whether scalar Lie point symmetries are variational for a large class of theories, simply by considering the actions of the symmetries on the kinetic terms. This criterion is independent of the specific parameter conditions associated with the often numerous potentials that realize the same symmetry algebra. The other two results, Proposition 2.5 and Corollary 1, were not strictly necessary for the present analysis, but they may substantially simplify the computations of variational, scalar Lie point symmetries of other, larger models of the type NHDM+KS. This is done by reducing the spacetime dimension, and hence the number of spacetime variables and gauge fields, while obtaining exactly the same results for scalar variational symmetries. Lie symmetry analysis detects only continuous (Lie) symmetries. Nevertheless, any discrete symmetries not detected by Lie’s method may be identified by considering the automorphism groups of the Lie symmetry algebras computed through Lie symmetry analysis. In this way, the set of possible maximal, faithfully acting variational symmetry groups of a generic model can be determined. A systematic study of such automorphisms and their role in identifying the discrete symmetries of large quantum-physical models is left for future work.
@@ Acknowledgments
The most general NHDM+KS Lagrangian in $d$ spacetime dimensions is given by (2.53). We can expand the Lagrangian (2.53) in purely scalar terms and terms linear and quadratic in the gauge fields, as follows:
@@
represent the gauge bosons and where e.g. $\mu\in\{0,1,2,3\}$ for $\mathcal{L}_{4}$ and $\mu=0$ for $\mathcal{L}_{1}$. Moreover, $T_{\text{GB}}$ is the sum of the kinetic terms of the gauge bosons given by (2.54), which is irrelevant for the proof of Proposition 2.5, because we only consider scalar symmetries. The relevant 1-prolongation of the scalar infinitesimal generator for the $d$-dimensional case is given by
@@
First we consider strict variational symmetries, that is, symmetries with $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$ (no sum over $d$), cf. (2.23) with $\xi=0$.
We denote $\partial_{\phi_{i}}\Phi_{j}\equiv\Phi_{j,i}$, and similar for Hermitian conjugated fields. Then e.g. $\Phi_{1,4}^{\dagger}=(0,-i)$. The effect of $\operatorname{pr}^{(1)}X_{d}$ of (A.3) on $\mathcal{L}_{d}$, given by (A), can be written as
@@
where $0\leq\mu\leq d-1$ and $1\leq K$. Assume that $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}=0$ holds, we would like to show that then $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$ holds, with the same $\eta$’s. This means that a scalar symmetry of the $d=1$ theory also is a scalar symmetry of the theory with arbitrary dimension $d$. Now, $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$, because all other terms are proportional to gauge bosons $G^{\mu}$ or derivatives $\partial\phi_{j}/\partial x_{\mu}$ (which includes terms $\partial_{\mu}s_{m}$), and hence cannot cancel $\eta_{i}{\partial_{\phi_{i}}V(\Phi,s)}{}$. This implies that $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$ in $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ as well, because it is the same expression when suppressing spacetime variables. The other terms all depend on $\mu$, and since they cancel for $\mu=0$ in $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}$, they also have to cancel for each $\mu$ in $\operatorname{pr}^{(1)}X_{4}\mathcal{L}_{4}$, because the structure of the terms is the same for all $\mu$.
@@
Conversely, assume $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$, that is, we have a symmetry of the theory of some fixed but arbitrary dimension $d>1$. Then, (A) holds where $\mu$ this time is a sum from 0 to d1d-1. We will show that then $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}=0$ for the same $\eta$’s, which means that the $d=1$ theory has the same symmetry. Again, because all other terms are proportional to gauge bosons or derivatives, $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$, and the same is true in the $d=1$ theory. Moreover, the other terms have to cancel for each $\mu$ individually, because terms involving gauge bosons $G^{\mu}$ cannot be canceled by terms involving gauge bosons $G^{\nu}$ for $\nu\neq\mu$, because they are independent fields. Moreover, the term $\frac{\partial\phi_{j}}{\partial{x_{\mu}}}\frac{\partial\eta_{i}}{\partial\phi_{j}}\frac{\partial\phi_{i}}{\partial x^{\mu}}$ for a specific, fixed $\mu$ cannot be canceled by a term of the same type with a fixed spacetime index $\nu\neq\mu$, because $\eta_{i}$ does not involve spacetime derivatives of fields. No other terms can cancel these terms either. Therefore, terms containing $\mu$ in (A) cancel for each choice of $\mu$ separately, and hence they will cancel for $\mu=0$ in the $d=1$ theory as well, and the symmetry for the $d>1$ theory is also a symmetry for the $d=1$ theory. Thus, the scalar, SVSs for $\mathcal{L}_{d}$ and $\mathcal{L}_{1}$ are the same. Hence, they are the same for $\mathcal{L}_{d}$ for all $d\in\mathbb{N}$.
@@
We now turn to the case where $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=\operatorname{Div}\beta\equiv d_{\mu}\beta^{\mu}$ for some local vector fields $\beta^{\mu}(x,\phi,\phi_{,\alpha},\phi_{,\alpha\beta})$, which are not identically zero. The first-order infinitesimal variation $\delta\mathcal{L}_{d}=\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ is a divergence if and only if the Euler operator $E$ applied to $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ is zero (cf. (2.28)), which means that
@@
for all $1\leq j\leq(4N+K+4d)$, where the dependent variables $y^{j}$ are the $4N+K$ Higgs field components and the $4d$ gauge field components, and where $d$ again is the number of independent variables (the dimension). First, assume $E(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$, i.e. we have a divergence symmetry in the $d=1$ theory. The term $\eta_{i}{\partial_{\phi_{i}}V(\Phi,s)}{}$ depends only on undifferentiated scalars; hence, if $E_{j}$ corresponds to a Higgs field $y^{j}=\phi_{j}$, then $E_{j}(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$ implies $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$. This is because the term $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))$ cannot be canceled by any other term in $E_{j}(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})$ because they are all proportional to either terms of the type $G_{\mu}$, $\partial_{\mu}\phi_{k}$ or a $\partial_{\mu}\partial^{\mu}\phi_{k}$ (all with $\mu=0$), in contrast to the term $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))$. Hence, $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$ will also be true in the $d>1$ theory, as these terms are the same for all $d$, suppressing spacetime variables. All the remaining terms of (A) have obtained two factors of the type $G_{\mu}$ or $\partial_{\mu}\phi_{k}$ for $\mu=0$, and regardless of which $E_{i}$ we apply, at least one such factor survives in each non-annihilated term; therefore, all these terms are tagged with the index $\mu=0$. Because the $d>1$ dimensional theory has $d$ copies of these terms, where each copy is tagged with an index $\mu$, $0\leq\mu\leq d-1$, each of the $d$ copies has to vanish separately because they are structurally the same, only distinguished by the index $\mu$. Thus, $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=0$.
@@
Conversely, assume $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=0$, that is, we have a divergence symmetry in $d>1$ spacetime dimensions. By the same reason as before the terms $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$ in the $d>1$ theory, and hence the same is true in the $d=1$ theory, since those terms are the same for all choices of $d$. The remaining terms of $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})$ can be grouped into $d$ distinct, non-overlapping categories determined by the value of $\mu$ in factors of the types $G_{\mu}$, $\partial_{\mu}\phi_{k}$ or $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$. The terms in each category must be canceled separately, since all terms in one category are linearly independent of any other term in another category. [^13] Note that the conditions for variational symmetries, including divergence symmetries, do not presuppose that the Euler-Lagrange equations of the Lagrangian hold. Hence, there is no relation between $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$ and $\partial_{\check{\nu}}\partial^{\check{\nu}}\phi_{k}$ for $\check{\mu}\neq\check{\nu}$. Hence the terms in the category $\mu=0$ cancel, and hence $E(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$. Therefore, the scalar symmetries of $\mathcal{L}_{d}$ are the same for all $d$, because $\mathcal{L}_{d}$ has the same scalar symmetries as $\mathcal{L}_{1}$ for all $d$. Finally, we note that the arguments hold equally well if $K=0$, that is, we have a pure NHDM. ∎
@@
We do not expect Proposition 2.5 to hold in general for non-variational symmetries, since the kinetic terms of the gauge bosons consist of cross terms of different spacetime derivatives, and these terms will now matter, because a symmetry of the Euler-Lagrange equations here implies that
@@
and the kinetic terms of the gauge bosons will be involved in the ”on-shell” condition $E(\mathcal{L}_{d})=0$, while this was not the case for the ”off-shell” (i.e., variational), scalar symmetries in Proposition 2.5. In contrast, the condition for a variational symmetry here is
@@
with no assumption on the validity of the Euler-Lagrange equations. Moreover, for $d=1$ the kinetic terms of the gauge bosons are zero because of the antisymmetry of the spacetime indices, whereas they are non-zero for $d>1$, which induces an asymmetry between the $d=1$ and $d>1$ cases in the on-shell condition $E(\mathcal{L}_{d})=0$ for the two cases. Finally, the Euler-Lagrange equations imply dependencies between derivatives such as $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$ and $\partial_{\check{\nu}}\partial^{\check{\nu}}\phi_{k}$ for $\check{\mu}\neq\check{\nu}$, in contrast to the off-shell situation in Proposition 2.5 (cf. footnote 13).